# Kernelized Concept Erasure

### v5

@@ 3.1 Background: Kernel Methods
Kernel methods are based on reproducing kernel Hilbert spaces (RKHS). Without going into the technical details, an RKHS is a space of “nice” functions equipped with a kernel (Yosida, 2012). A kernel $\kappa(\cdot,\cdot)\geq 0$ is a similarity measure that generalizes a positive definite matrix. When we have a kernel over an RKHS, then the kernel corresponds to the dot product of a feature map, i.e., $\kappa(\mathbf{x},\mathbf{y})=\boldsymbol{\Phi}(\mathbf{x})^{\top}\boldsymbol{\Phi}(\mathbf{y})$. This insight gives us a natural manner to construct kernels. For instance, the linear kernel $\kappa(\mathbf{x},\mathbf{y})=\mathbf{x}^{\top}\mathbf{y}$ corresponds to the standard dot product in Euclidean space. The degree-2 polynomial kernel $\kappa(\mathbf{x},\mathbf{y})={(\gamma\mathbf{x}^{\top}\mathbf{y}+\alpha)}^{2}$ corresponds to a dot product in a six-dimensional feature space if $\mathbf{x},\mathbf{y}\in\mathbb{R}^{2}$.
Kernels are more general than positive definite matrices in that they can exist in infinite-dimensional spaces. For example, the Gaussian kernel $\exp\left(-\gamma||\mathbf{x}-\mathbf{y}||_{2}^{2}\right)$ is infinite-dimensional. [^3] One way to see this is to show a Gram matrix of any size $N$ will always have full rank. However, for any finite set of points $\{\mathbf{x}_{1},\dots,\mathbf{x}_{N}\}$, we can construct a Gram matrix $K\in\mathbb{R}^{N\times N}$ where $K_{nm}=\kappa(\mathbf{x}_{n},\mathbf{x}_{m})$ encodes the similarity between $\mathbf{x}_{n}$ and $\mathbf{x}_{m}$. The matrix $K$ is guaranteed to be positive definite. Kernels are useful because they allow us to implicitly learn functions in a potentially infinite-dimensional RKHS without materializing that space.
3.2 A Kernelized Minimax Game
Inspection of the linear adversarial game in Eq. 1 reveals that both the adversary and the predictor interact with the input only via an inner product. Thus, the game can be kernelized by replacing the inner product with a kernel operation. We establish this kernelization by first proving the following representer theorem-like lemma, which shows that $\boldsymbol{w},{\boldsymbol{\theta}}$ can be written in terms of spans of the projected training set.
@@ 3.2 A Kernelized Minimax Game
(Minimax Game Representer Theorem) Let $\mathcal{H}$ be a reproducing kernel Hilbert space with canonical feature map $\boldsymbol{\Phi}:\mathbb{R}^{D}\rightarrow\mathcal{H}$, i.e., $\boldsymbol{\Phi}(\mathbf{x})=\kappa(\mathbf{x},\cdot)$. Consider the game:
$\max_{\boldsymbol{w}\in\mathcal{H}}\min_{{\boldsymbol{\theta}}\in\mathcal{H}}\sum_{n=1}^{N}\ell\left(y_{n},\,\langle{\boldsymbol{\theta}},\mathrm{P}_{\boldsymbol{w}}^{\perp}\,\boldsymbol{\Phi}(\mathbf{x}_{n})\rangle\right)$ (5)
[…]
@@
Having expressed $\boldsymbol{w},{\boldsymbol{\theta}}$ as a function of the training data, we can proceed to proving the kernelization of the adversarial game.
@@
[…]
$-\frac{\boldsymbol{\alpha}^{\top}K^{(m)}(\mathbf{z})\boldsymbol{\alpha}}{\boldsymbol{\alpha}^{\top}K\boldsymbol{\alpha}}\bigg{)}$ (6)
where $K^{(m)}_{ij}(\mathbf{z})\mathrel{\stackrel{{\scriptstyle\textnormal{def}}}{{=}}}\kappa(\mathbf{x}_{i},\mathbf{z})\kappa(\mathbf{x}_{m},\mathbf{x}_{j})$.
@@
$\min_{\boldsymbol{\beta}\in\mathbb{R}^{N}}\,\max_{\boldsymbol{\alpha}\in\mathbb{R}^{N}}\,\sum_{n=1}^{N}\ell\Bigg{(}y_{n},\,$ $\sum_{m=1}^{N}\beta_{m}\bigg{(}\kappa(\mathbf{x}_{m},\mathbf{z}_{n})$
$-\frac{\boldsymbol{\alpha}^{\top}K^{(m,n)}\boldsymbol{\alpha}}{\boldsymbol{\alpha}^{\top}K\boldsymbol{\alpha}}\bigg{)}\Bigg{)}$ (7)
where we define $K^{(m,n)}_{ij}(\mathbf{z})\mathrel{\stackrel{{\scriptstyle\textnormal{def}}}{{=}}}\kappa(\mathbf{x}_{i},\mathbf{z}_{n})\kappa(\mathbf{x}_{m},\mathbf{x}_{j})$. In contrast to Eq. 5, all computations in § 3.2 are in Euclidean space.
@@
The reproducing kernel Hilbert space game Eq. 5 attains the same local optima as § 3.2.
@@
The objective in § 3.2 can be computed in $\mathcal{O}(N^{4})$ time.
@@
Assuming that $\kappa(\cdot,\cdot)$ may be computed in $\mathcal{O}(1)$, computing $K^{(m,n)}$ takes $\mathcal{O}(N^{2})$ time. We have to do $\mathcal{O}(N^{2})$ such computations, which results in $\mathcal{O}(N^{4})$ time. Note that $K$ may be pre-computed once in $\mathcal{O}(N^{2})$ time. Thus, $K^{(m,n)}$ is the bottleneck, so the whole algorithm takes $\mathcal{O}(N^{4})$ time.∎
There are two problems with naïvely using the formulation in § 3.2. First, and similarly to Eq. 1, the problem is not convex–concave due to the optimization over $\boldsymbol{\alpha}$, which implicitly defines an orthogonal projection matrix of rank 1. Second, the evaluation time of § 3.2 is $\mathcal{O}\left(N^{4}\right)$, as argued in Proposition 1, which makes using a training set larger than a few hundred examples infeasible. We solve both of these computational issues with the Nyström approximation.
@@ 3.3 The Nyström Approximation
3.3.1 Convexifying the Objective
Consider the Gram matrix $K\in\mathbb{R}^{N\times N}$ where $K_{nm}=\kappa(\mathbf{x}_{n},\mathbf{x}_{m})$, $\mathbf{x}_{n}$ is the $n^{\text{th}}$ training representation, and $\mathbf{x}_{m}$ is the $m^{\text{th}}$ training representation. We start with the eigendecomposition of $K=U\Sigma U^{\top}=U\sqrt{\Sigma}\sqrt{\Sigma}U^{\top}$, which can be computed in $\mathcal{O}\left(N^{3}\right)$ time (Golub and Van Loan, 2013). We are justified in taking the square root of $\Sigma$ because $K$ is necessarily positive definite. Now, we define an approximate feature map for observations $\mathbf{x}_{n}$ in the training data using the eigenvalues and vectors:
$\boldsymbol{\widetilde{\Phi}}(\mathbf{x}_{n})\mathrel{\stackrel{{\scriptstyle\textnormal{def}}}{{=}}}\left(U\sqrt{\Sigma}\right)_{n}$ (8)
To compute the features for representations $\mathbf{x}$ not in the training data, we use the following:
$\widetilde{\boldsymbol{\Phi}}\left(\mathbf{x}\right)\mathrel{\stackrel{{\scriptstyle\textnormal{def}}}{{=}}}\sum_{n=1}^{N}\kappa(\mathbf{x},\mathbf{x}_{n})\widetilde{\boldsymbol{\Phi}}(\mathbf{x}_{n})$ (9)

### v6

@@ 3.1 Background: Kernel Methods
Kernel methods are based on reproducing kernel Hilbert spaces (RKHS). Without going into the technical details, an RKHS is a space of “nice” functions equipped with a kernel (Yosida, 2012). A kernel ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\kappa}(\cdot,\cdot)\geq 0$ is a similarity measure that generalizes the notion of a positive definite matrix. When we have a kernel over an RKHS, the kernel corresponds to the dot product of a feature map, i.e., ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\kappa}(\boldsymbol{x},\boldsymbol{y})=\boldsymbol{\Phi}(\boldsymbol{x})^{\top}\boldsymbol{\Phi}(\boldsymbol{y})$, for feature map $\boldsymbol{\Phi}$. This insight gives us a natural manner to construct kernels. For instance, the linear kernel ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\kappa}(\boldsymbol{x},\boldsymbol{y})=\boldsymbol{x}^{\top}\boldsymbol{y}$ corresponds to the standard dot product in Euclidean space. The degree-2 polynomial kernel ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\kappa}(\boldsymbol{x},\boldsymbol{y})={(\gamma\boldsymbol{x}^{\top}\boldsymbol{y}+\alpha)}^{2}$ corresponds to a dot product in a six-dimensional feature space if $\boldsymbol{x},\boldsymbol{y}\in\mathbb{R}^{2}$ where $\gamma\in\mathbb{R}$ is a hyperparameter.
Kernels are more general than positive definite matrices in that they can exist in infinite-dimensional spaces. For example, the Gaussian kernel $\exp\left(-\gamma||\boldsymbol{x}-\boldsymbol{y}||_{2}^{2}\right)$ is infinite-dimensional, where, again, $\gamma\geq 0$ is a hyperparameter. [^3] One way to see this is to show a Gram matrix of any size $N$ will always have full rank. However, for any finite set of points $\{\boldsymbol{x}_{1},\dots,\boldsymbol{x}_{N}\}$, we can construct a Gram matrix ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\boldsymbol{K}}\in\mathbb{R}^{N\times N}$ where ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\boldsymbol{K}^{(m,n)}}={\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\kappa}(\boldsymbol{x}_{n},\boldsymbol{x}_{m})$ encodes the similarity between $\boldsymbol{x}_{n}$ and $\boldsymbol{x}_{m}$. The matrix ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\boldsymbol{K}}$ is guaranteed to be positive definite. Kernels are useful because they allow us to implicitly learn functions in a potentially infinite-dimensional RKHS without materializing that space.
3.2 A Kernelized Maximin Game
Inspection of the maximin game in Eq. 1 reveals that both the adversary and the predictor interact with the input only via an inner product. Thus, the game can be kernelized by replacing the inner product with a kernel operation. We establish this kernelization by first proving the following representer theorem-like lemma, which shows that $\boldsymbol{w}$ and ${\boldsymbol{\theta}}$, now considered to live an RKHS $\mathcal{H}$, can be written in terms of spans of the projected training set.
@@ 3.2 A Kernelized Maximin Game
(Maximin Game Representer Theorem) Let $\mathcal{H}$ be a reproducing kernel Hilbert space with canonical feature map $\boldsymbol{\Phi}\colon\mathbb{R}^{D}\rightarrow\mathcal{H}$, i.e., $\boldsymbol{\Phi}(\mathbf{x})={\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\kappa}(\mathbf{x},\cdot)$. Consider the following maximin game:
$\max_{\boldsymbol{w}\in\mathcal{H}}\min_{{\boldsymbol{\theta}}\in\mathcal{H}}\sum_{n=1}^{N}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\ell}\left(y_{n},\,\langle{\boldsymbol{\theta}},\mathrm{P}_{\boldsymbol{w}}^{\perp}\,\boldsymbol{\Phi}(\boldsymbol{x}_{n})\rangle\right).$ (5)
where $\mathrm{P}_{\boldsymbol{w}}^{\perp}\colon\mathcal{H}\rightarrow\mathcal{H}$ is the operator that projects onto the orthogonal complement of $\boldsymbol{w}$. For every attained local optimum $({\boldsymbol{\theta}}^{\star}$, $\boldsymbol{w}^{\star})$ of Eq. 5, there is another local optimum $({\boldsymbol{\theta}}_{U}^{\star}$, $\boldsymbol{w}_{U}^{\star})$ with the same value as $({\boldsymbol{\theta}}^{\star}$, $\boldsymbol{w}^{\star})$ in $U\mathrel{\stackrel{{\scriptstyle\textnormal{def}}}{{=}}}\texttt{span}\Big{\{}\boldsymbol{\Phi}(\mathbf{x}_{1}),\ldots,\boldsymbol{\Phi}(\mathbf{x}_{N})\Big{\}}$, the span of the training data. [^4] Traditionally, a representer theorem has an extra regularization term on the function estimator that ensures the optimum is attained and that it is global (Kimeldorf and Wahba, 1970). We do not employ such a regularizer and, thus, have a weaker result.
@@
Having expressed $(\boldsymbol{w},{\boldsymbol{\theta}})$ as a function of the training data, we can proceed to proving the kernelization of the adversarial game.
@@
Let $\mathcal{H}$ be a reproducing kernel Hilbert space with canonical feature map $\boldsymbol{\Phi}$, and let $\boldsymbol{\Phi}(\boldsymbol{z})$ be a point in $\mathcal{H}$. Next, let $\boldsymbol{w}=\sum_{n=1}^{N}\alpha_{n}\boldsymbol{\Phi}(\boldsymbol{x}_{n})$ and ${\boldsymbol{\theta}}=\sum_{n=1}^{N}\beta_{n}\boldsymbol{\Phi}(\boldsymbol{x}_{n})$ be points in the reproducing kernel Hilbert space. Now, let ${\boldsymbol{\Phi}_{\text{proj}}(\boldsymbol{z})}$ be the orthogonal projection of $\boldsymbol{\Phi}(\boldsymbol{z})$ onto the orthogonal complement of the subspace spanned by $\boldsymbol{w}$. Then, we have
$\langle{\boldsymbol{\theta}},{\boldsymbol{\Phi}_{\text{proj}}(\boldsymbol{z})}\rangle=\sum_{m=1}^{N}$ $\beta_{m}\bigg{(}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\kappa}(\boldsymbol{x}_{m},\boldsymbol{z})$ (6)
$-\frac{\boldsymbol{\alpha}^{\top}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\boldsymbol{K}^{(m)}}(\boldsymbol{z})\boldsymbol{\alpha}}{\boldsymbol{\alpha}^{\top}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\boldsymbol{K}}\boldsymbol{\alpha}}\bigg{)},$
where ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\boldsymbol{K}}^{(m)}_{ij}(\boldsymbol{z})\mathrel{\stackrel{{\scriptstyle\textnormal{def}}}{{=}}}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\kappa}(\boldsymbol{x}_{i},\boldsymbol{z}){\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\kappa}(\boldsymbol{x}_{m},\boldsymbol{x}_{j})$.
@@
$\max_{\boldsymbol{\alpha}\in\mathbb{R}^{N}}\,\min_{\boldsymbol{\beta}\in\mathbb{R}^{N}}\,\sum_{n=1}^{N}$ ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\ell}\Bigg{(}y_{n},\sum_{m=1}^{N}\beta_{m}\bigg{(}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\kappa}(\boldsymbol{x}_{m},\boldsymbol{z}_{n})$ (7)
$-\frac{\boldsymbol{\alpha}^{\top}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\boldsymbol{K}^{(m,n)}}\boldsymbol{\alpha}}{\boldsymbol{\alpha}^{\top}K\boldsymbol{\alpha}}\bigg{)}\Bigg{)},$
where we define ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\boldsymbol{K}}^{(m,n)}_{ij}(\boldsymbol{z})\mathrel{\stackrel{{\scriptstyle\textnormal{def}}}{{=}}}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\kappa}(\boldsymbol{x}_{i},\boldsymbol{z}_{n}){\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\kappa}(\boldsymbol{x}_{m},\boldsymbol{x}_{j})$. In contrast to Eq. 5, all computations in Eq. 7 are in Euclidean space.
@@
The reproducing kernel Hilbert space game Eq. 5 attains the same local optima as Eq. 7.
@@
The objective in Eq. 7 can be computed in $\mathcal{O}(N^{4})$ time.
@@
Assuming that ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\kappa}(\cdot,\cdot)$ may be computed in $\mathcal{O}(1)$, computing ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\boldsymbol{K}^{(m,n)}}$ takes $\mathcal{O}(N^{2})$ time. We have to do $\mathcal{O}(N^{2})$ such computations, which results in $\mathcal{O}(N^{4})$ time. Note that $K$ may be pre-computed once in $\mathcal{O}(N^{2})$ time. Thus, ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\boldsymbol{K}^{(m,n)}}$ is the bottleneck, so the whole algorithm takes $\mathcal{O}(N^{4})$ time.∎
There are two problems with naïvely using the formulation in Eq. 7. First, and similarly to Eq. 1, directly optimizing Eq. 7 is difficult because the parameters $\boldsymbol{\alpha}$ implicitly define an orthogonal projection matrix of rank 1 and, thus, we face the sample problems discussed in Ravfogel et al. (2022). Second, the evaluation time of Eq. 7 is $\mathcal{O}\left(N^{4}\right)$, as argued in Proposition 1, which makes using a training set larger than a few hundred examples infeasible. We solve both of these computational issues with the Nyström approximation.
@@ 3.3 The Nyström Approximation
3.3.1 Relaxing the Objective
Consider the Gram matrix ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\boldsymbol{K}}\in\mathbb{R}^{N\times N}$ where ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\boldsymbol{K}^{(m,n)}}={\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\kappa}(\boldsymbol{x}_{m},\boldsymbol{x}_{n})$, $\boldsymbol{x}_{m}$ is the $m^{\text{th}}$ training representation, and $\boldsymbol{x}_{n}$ is the $n^{\text{th}}$ training representation. We start with the eigendecomposition of ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\boldsymbol{K}}={\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\boldsymbol{U}}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\boldsymbol{\Sigma}}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\boldsymbol{U}}^{\top}={\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\boldsymbol{U}}\sqrt{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\boldsymbol{\Sigma}}}\sqrt{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\boldsymbol{\Sigma}}}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\boldsymbol{U}}^{\top}$, which can be computed in $\mathcal{O}\left(N^{3}\right)$ time (Golub and Van Loan, 2013). We are justified in taking the square root of ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\boldsymbol{\Sigma}}$ because ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\boldsymbol{K}}$ is necessarily positive definite. Now, we define an approximate feature map for observations $\boldsymbol{x}_{n}$ in the training data using the eigenvalues and vectors:
$\boldsymbol{\widetilde{\Phi}}(\boldsymbol{x}_{n})\mathrel{\stackrel{{\scriptstyle\textnormal{def}}}{{=}}}\left({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\boldsymbol{U}}\sqrt{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\boldsymbol{\Sigma}}}\right)_{n}$ (8)
To compute the features for representations $\boldsymbol{x}$ not in the training data, we use the following:
$\widetilde{\boldsymbol{\Phi}}\left(\boldsymbol{x}\right)\mathrel{\stackrel{{\scriptstyle\textnormal{def}}}{{=}}}\sum_{n=1}^{N}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\kappa}(\boldsymbol{x},\boldsymbol{x}_{n})\widetilde{\boldsymbol{\Phi}}(\boldsymbol{x}_{n})$ (9)